APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday August 28, 2009 (MJD 55072) ---=== OBSERVING TEAM ===--- Alaina Shelden Steph Snedden Robert Lupton (PU) Paul Harding (CWRU) Kyle Dawson (U of U) Demitri Muna (NYU) David Schlegel (LBL) ---=== OBSERVING PLAN ===--- More guider software development. ---=== OBSERVING SUMMARY ===--- Lots of pounding, pushing and pulling on the guider software. Good progress was made on getting the guider loop to work. A scaleChange command was created. Most needed commands now exist from the command line. The spectro autofill problem is still with us, unfortunately. See the log body. We force-filled the spectrographs at the end of the night, but left the autofill off. SOMEONE WILL NEED TO CHECK/FILL THE SP DEWARS DURING THE DAY. The Mac Mini froze at the end of the night. At the moment, it is back up. Jon, please read the details in the log body. ---=== OBSERVING LOG ===--- Afternoon: ---------- Afternoon checkout was okay, except the autofill problem is not resolved, apparently. The intermediate spectro dewar (sp2 we think, but we only noticed this from the control room, so can't be sure) stopped and restarted the fill sequence before the dewar was anywhere close to being full. We stopped the fill so we could manually finish it after opening. Did not check disk space, since we are working on hub25m, not sdsshost2. Night: ----- 01:37Z - Opened enclosure. Finished the manual fill of the spectrographs. The dewars must have been close to empty, since it took between 60 and 70 lbs to fill. 02:23Z - Waiting for the first BOSS field to rise to an acceptable altitude. 03:10Z - Cartridge 11, plate 3520 - BOSS One goal was to implement/troubleshoot the guider offset loop. This was successful for the most part. The rotator offsets need a little more work. Paul sez "we flipped about 50 different signs". Also, Robert implemented a scaleChange command. Many of the commands we need to use for operations are now available from the command line. 11:09Z - Cartridge 10, plate 3523 - BOSS Fine-tuned the fiber angles for this plate. 11:21Z - The brand new Mac Mini choked (froze) and would not reboot. It felt hot. We turned it off. Alaina found a dcument on how to reset the power management unit, which we were forced to do, since it would not reboot. The clock and other settings may need to be reset now. The fan came on as soon as we reset the PMU, and we don't recall hearing the fan at all before. 12:02Z - Closed enclosure. 12:13Z - Force-filled spectro intermediate dewars. 12:13Z - endNight finished. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:36Z 61 34 12 83 (E) 599 - mostly clear 00:08Z 61 31 16 71 (ENE) 571 - ' 00:40Z 60 26 20 74 (ENE) 523 - ' 01:13Z 59 31 26 64 (ENE) 565 - ' 01:45Z 58 30 26 66 (ENE) 617 - ' 02:17Z 57 31 19 82 (E) 601 - ' 02:49Z 57 31 16 74 (ENE) 650 - ' 11:44Z 54 34 13 78 (ENE) 1244 - ' 12:16Z 53 34 6 85 (E) 1219 - ' ---=== TELESCOPE STATUS ===--- 00:45Z - Started ventilation 01:37Z - Opened enclosure 12:02Z - Closed enclosure Telescope stowed at 30 deg alt with BOSS cartridge 10 mounted. Counterweights at 230 Instrument lift/latches in manual mode Spectro autofill in MANUAL MODE 12:08Z - spectro dewar: 190 lb; 15 psi and still filling 12:08Z - imager dewar: 131 lb; 20 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 Watcher: v2_36_0 MCP: NOSVN TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.11.0 2009-06-24 sdssProcedures: v2_44 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.6920 -0.6940 Axial B -0.0440 -0.0410 Axial C -0.3740 -0.3710 Trans D 11.6090 11.6170 Lateral E 10.3505 10.3505 Lateral F 11.9253 11.9253 GALILS Commanded: 9900. 400. 5850. 1200. 2950. 2750. Actual: 9917. 401. 5832. 1205. 2954. 2757. SETMIR VALUES PriDesOrient: 0.00 -3.30 5.10 278.08 178.06 PriOrient: 0.00 -3.04 5.26 277.70 178.07 SECONDARY: ---------- Focus: 0.00 Air Temp.: 15.3 Alt.: 30.002054 MIGS TONIGHT NOMINAL Axial A 18.6160 18.6150 Axial B 15.6140 15.6120 Axial C 13.8180 13.8170 Trans D -11.7760 -11.7780 GALILS Commanded: 1597652. 1566188. 1588090. -7600. -12400. Actual: 1597721. 1566188. 1588095. -7597. -12380. SETMIR VALUES SecDesOrient: 1257.00 -4.00 -9.00 10.00 234.42 SecOrient: 1256.98 -4.01 -8.98 10.20 234.69 ---=== PROBLEMS IN DETAIL ===---